# Gravitational radiation reaction around a static black hole surrounded by a Dehnen type dark matter halo

### v1

@@ I Introduction
The interplay between black holes and DM has proved to be a significant field of study in theoretical astrophysics, especially in understanding the dynamics of gravitational systems Boshkayev et al. (2020, 2021); Capozziello et al. (2025). Black holes play a crucial role in the formation and evolution of galaxies. On the other hand, DM, which constitutes about $27\%$ of the universe’s total mass-energy content, adds further complication to these gravitational systems and in cosmology Dunsby et al. (2016). In this context, the study of gravitational radiation reaction around black holes immersed in a DM halo offers deep insights into how these entities influence each other’s behavior and the resulting GW emitted.
@@
The presence of extended DM halos surrounding galaxies is strongly supported by various lines of evidence, including galaxy rotation curves, gravitational lensing analyses, and measurements of cosmic microwave background anisotropies. Among the models used to describe such halos, the Navarro–Frenk–White (NFW) profile Navarro et al. (1996), derived from cold DM simulations, is widely employed and is characterized by its steep inner density cusp. Nevertheless, the NFW profile has limited adaptability for representing alternative halo morphologies, such as cored density profiles. On the other hand, the Dehnen profile Dehnen (1993a) offers a more general double power-law form, encompassing the Jaffe and Hernquist models as specific cases and featuring an adjustable inner slope. This versatility makes it especially suitable for capturing both cuspy and cored DM configurations, which are relevant for modeling the environments of black holes Stegmann et al. (2020a); Hamil et al. (2025).
@@
Recent advances in astrophysical research have increasingly highlighted the relevance of the Dehnen DM halo model. Since its introduction as a spherically symmetric density profile, the Dehnen model has gained widespread application in galactic dynamics and black hole astrophysics Dehnen (1993a). Initially employed to characterize the nuclear bulges of elliptical galaxies, it has more recently been used to construct black hole models embedded within Dehnen-profiled DM halos, revealing the impact of DM on the gravitational environment of black holes Pantig and Övgün (2022). Further studies have derived black hole solutions for specific Dehnen profiles, systematically analyzing their thermodynamic behavior and strong-field gravitational signatures Gohain et al. (2024a). Investigations have also demonstrated that the gravitational lensing properties and orbital dynamics predicted by this model align well with general relativity, making it a preferred theory for explaining the peripheral environment of galaxies Ali et al. (2025). Collectively, these results establish the Dehnen profile as a powerful framework for probing DM–black hole interactions and the structure of galactic environments.
@@
In addition, GWs have become of large interest in gravitational physics since its detection by LIGO/Virgo collaboration in 2015 Abbott et al. (2016). One of the main sources of GW is two body systems in which a stellar-mass compact object (SCO) rotates a supermassive black hole (SMBH). In such extreme mass ratio inspirals (EMRIs), low frequency GW would be produced which could be observed in future space-based observatories like LISA Arun et al. (2022).
@@
In this paper, we are interested in EMRIs immersed in a DM halo whose density distribution is given by Dehnen type profile Dehnen (1993b). Dehnen distribution has been used, by several authors, to model the DM profile of dwarf galaxies Stegmann et al. (2020b); Inoue (2017); Errani et al. (2018). The traditional belief is that these galaxies rarely host a SMBH at their centers. However, some recent observations, including that of Leo I galaxy, show that SMBHs might be present at the center of some dwarf galaxies Bustamante-Rosell et al. (2021) (see also Ref. Reines (2022)). This has motivated study of black holes embedded in Dehnen type DM halo Pantig and Övgün (2022); Al-Badawi et al. (2025); Gohain et al. (2024b); Al-Badawi and Shaymatov (2025a, b). As the SCO spiral into the SMBH, it emits GWs. Further, the DM halo exerts a gravitational drag force on SCO which is known as dynamical friction. On the other hand, SCO accretes DM as it is moving inside the halo. We are interested in studying the evolution of the orbital parameter of SCO as well as its energy and angular momentum due to GW radiation reaction, dynamical friction, and the accretion. Various black hole solution surrounded by DM halo have been found in literature Cardoso et al. (2021, 2022a, 2022b); Figueiredo et al. (2023) and the evolution of orbits of small compact objects inspiralling around these black holes has been investigated Dai et al. (2022); Zhang et al. (2024). These studies could help us find out about the existence and properties of DM.
@@
The paper is structured as follows. In Sec. II, we describe the main features of dark matter density profile, computing the corresponding mass associated with dark matter. In Sec. III, the bound orbits of the underlying metric is explored. In Sec. IV, the gravitational waveforms are studied, and finally in Sect. V, we sketch our conclusions and perspectives.
II Dark matter density profile and metric function
@@
$M_{D}=4\pi\int\limits_{0}^{r}\rho\left(r^{\prime}\right)r^{\prime 2}dr^{\prime}.$ (1)
@@
$\rho=\rho_{s}\left(\frac{r}{r_{s}}\right)^{-\gamma}\left[\left(\frac{r}{r_{s}}\right)^{\alpha}+1\right]^{\frac{\gamma-\beta}{\alpha}},$ (2)
where $\rho_{s}$ and $r_{s}$ denote the central halo density and the halo core radius respectively, whereas $\gamma$ determines the specific variant of the profile. The values of $\gamma$ lies within $[0,3]$, e.g. $\gamma=3/2$ is used to fit the surface brightness profiles of elliptical galaxies which closely resembles the de Vaucouleurs $r^{1/4}$ profile Shakeshaft et al. (1974).
@@
Accordingly, Eq. (2) becomes
$\rho=\frac{\rho_{s}}{\left(\frac{r}{r_{s}}\right)^{5/2}\left(\frac{r}{r_{s}}+1\right)^{3/2}},$ (3)
and, so, plugging Eq. (3) in Eq. (1), one obtains
$M_{D}=4\pi\int\limits_{0}^{r}\frac{\rho_{s}r^{\prime 2}}{\left(\frac{r^{\prime}}{r_{s}}\right)^{5/2}\left(\frac{r^{\prime}}{r_{s}}+1\right)^{3/2}}dr^{\prime}=\frac{8\pi\rho_{s}r_{s}^{3}}{\sqrt{1+\frac{r_{s}}{r}}}.$ (4)
The static black hole metric in the Dehnen-$(1,4,5/2)$ DM halo is therefore given by  [^1] The authors of Ref. Al-Badawi et al. (2025) model DM using an anisotropic fluid stress-energy tensor. Similarly, an effective anisotropic fluid has been proposed in Refs. Cadoni et al. (2018a, b); Giusti et al. (2022) to unify dark energy and DM. The procedure to model, with an external energy-momentum tensor, induced by DM, a background metric appears a viable approximation to characterize with one metric only, two different spacetimes, i.e., an interior and an exterior configuration. Clearly, this limits the overall description, since it exhibits the presence of pressures violating the the cosmological cold DM hypothesis. However, it is possible to handle this technique at first approximation, albeit a more precise and complete scheme would require the Israel-Darmois junction conditions Chu and Tan (2022) or alternatives, see e.g. Luongo and Quevedo (2024, 2014).  Al-Badawi et al. (2025)
$ds^{2}=-f\left(r\right)dt^{2}+\frac{dr^{2}}{f\left(r\right)}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (5)
@@
$f\left(r\right)=1-\frac{2M}{r}-32\pi\rho_{s}r_{s}^{3}\sqrt{\frac{r+r_{s}}{r_{s}^{2}\,r}}\,.$ (6)
This choice, though restrictive compared to the more general ansatz $g_{tt}=-e^{\Phi(r)}f(r)$, which can describe spacetimes with anisotropic matter distributions, is consistent with the weak-field origin of the DM halo model adopted in Ref. Al-Badawi et al. (2025) and ensures compliance with standard energy conditions. Note that this geometry differs from the vacuum Schwarzschild black hole and the inclusion of DM has changed the horizon radius (see e.g. Ref. Xu et al. (2018)).
III Bound orbits of the corresponding black hole
@@ III.1 The geodesic equations
$\frac{dt}{d\tau}=\frac{E}{f(r)},$ (7)
$\frac{d\phi}{d\tau}=\frac{L}{r^{2}},$ (8)
$\left(\frac{dr}{d\tau}\right)^{2}+V(r;L)=E^{2},$ (9)
where $\tau$ is the particle’s proper time222 Since the analysis is restricted to the weak-field region where $f(r)\approx 1$, the difference between $t$ and proper time $\tau$ is negligible. This is a standard simplification for large-radius orbits in similar studies.; $E$ and $L$ are constants of motion, respectively, the orbital energy and angular momentum; the effective potential for radial motion is given by
$V(r;L)=f(r)\left(1+\frac{L}{r^{2}}\right).$ (10)
@@ III.2 Orbital parameters: $p$ and $e$
$\frac{\partial V}{\partial r}=\frac{\partial^{2}V}{\partial r^{2}}=0.$ (11)
@@
$L_{\rm bound}=\frac{2\sqrt{3}\left(M+8\pi\rho_{s}r_{s}^{3}\right)}{1-32\rho_{s}r_{s}^{2}}.$ (12)
@@
The roots of Eq. (13) can be obtained by defining Hughes (2024)
$R_{3}(r)=r^{3}+\mathcal{A}_{1}r^{2}+\mathcal{A}_{2}r+\mathcal{A}_{3}\;,$ (14)
@@
$\mathcal{A}_{1}$ $=$ $-\frac{2M+16\pi\rho_{s}r_{s}^{3}}{(1-E^{2}-32\pi\rho_{s}r_{s}^{2})}\;,$ (15)
$\mathcal{A}_{2}$ $=$ $\frac{L^{2}(1-32\pi\rho_{s}r_{s}^{2})}{(1-E^{2}-32\pi\rho_{s}r_{s}^{2})}\;,$ (16)
$\mathcal{A}_{3}$ $=$ $-\frac{2L^{2}(M+8\pi\rho_{s}r_{s}^{3})}{(1-E^{2}-32\pi\rho_{s}r_{s}^{2})}\;.$ (17)
@@
$\mathcal{Q}$ $=$ $\frac{1}{9}\left(\mathcal{A}_{1}^{2}-3\mathcal{A}_{2}\right)\;,\;$ (18)
$\mathcal{R}$ $=$ $\frac{1}{54}\left(2\mathcal{A}_{1}^{3}-9\mathcal{A}_{1}\mathcal{A}_{2}+27\mathcal{A}_{3}\right),$ (19)
@@
$\vartheta=\arccos\left(\mathcal{R}/\sqrt{\mathcal{Q}^{3}}\right)\;.$ (20)
@@
$r_{\rm 1}$ $=$ $-2\sqrt{\mathcal{Q}}\cos\left(\frac{\vartheta+2\pi}{3}\right)-\frac{\mathcal{A}_{1}}{3}\;,$ (21)
$r_{\rm 2}$ $=$ $-2\sqrt{\mathcal{Q}}\cos\left(\frac{\vartheta-2\pi}{3}\right)-\frac{\mathcal{A}_{1}}{3}\;,$ (22)
$r_{3}$ $=$ $-2\sqrt{\mathcal{Q}}\cos\left(\frac{\vartheta}{3}\right)-\frac{\mathcal{A}_{1}}{3}\;.$ (23)
For a SCO inspiraling into a SMBH surrounded by a DM halo, the geodesic motion of the SCO can be parameterized by the semimajor axis $a$ and eccentricity $e$ through Barack and Long (2022)
$r=\frac{a(1-e^{2})}{1+e\cos\chi},$ (24)
@@
For eccentric orbits, it is useful to relabel the outer two roots, which define the range of the smaller body’s motion: the outermost root named as the apoapsis, and the next root as the periapsis: $r_{1}=r_{a}$, $r_{2}=r_{p}$. From Eqs. (21) - (23), the third root can be written in terms of $p$ and $e$. Stable eccentric orbits occur when the inequality $r_{3}<r_{2}<r_{1}$ is satisfied, while stable circular orbits occur when $r_{1}=r_{2}>r_{3}$. The relations between $\left\{p,e\right\}$ and $\left\{E,L\right\}$ can be obtained as
$\epsilon$ $=$ $E/\mu=\sqrt{\frac{f(r_{a})f(r_{p})\left(r_{a}^{2}-r_{p}^{2}\right)}{r_{a}^{2}f(r_{p})-r_{p}^{2}f(r_{a})}},$ (25)
$\ell$ $=$ $L/\mu=\sqrt{\frac{r_{a}^{2}r_{p}^{2}\left(f(r_{a})-f(r_{p})\right)}{r_{a}^{2}f(r_{p})-r_{p}^{2}f(r_{a})}}.$ (26)
Combining Eq. (9) and Eqs. (24)-(26) yields
@@
$P$ $=$ $\int_{0}^{2\pi}\frac{dt}{d\chi}d\chi,$ (29)
$\Delta\phi$ $=$ $\int_{0}^{2\pi}\frac{d\phi}{d\chi}d\chi-2\pi.$ (30)
@@
EMRIs occurring within DM halos emit GWs and are affected by their ambient environment. The SCO’s motion is shaped by several forces, including GW backreaction, dynamical friction, and material accretion from the halo. If the SCO is a small black hole, it will draw in surrounding DM as it travels through the halo, a process that can be modeled using the Bondi-Hoyle accretion formula Macedo et al. (2013)
$\dot{\mu}=\frac{4\pi\mu^{2}\rho_{\text{DM}}}{(v^{2}+c_{\text{s}}^{2})^{3/2}},$ (31)
@@
$\left(\frac{dE}{dt}\right)_{\text{acc}}=\frac{\dot{\mu}}{\mu}E+\mu\dot{\epsilon}=\frac{\dot{\mu}}{\mu}\left(E-{\ell}\frac{dp}{d\ell}\frac{dE}{dp}\right).$ (32)
As the small black hole moves through a DM environment, it experiences gravitational interactions with DM particles within the halo. These interactions produce dynamical friction, which contributes to the loss of both energy and angular momentum. The Chandrasekhar formula for dynamical friction is expressed as Chandrasekhar (1943):
${\bf f}_{\text{DF}}=-\frac{4\pi\mu^{2}\rho_{\text{DM}}\ln\Lambda}{v^{3}}{\bf v}\bm{v},$ (33)
where the Coulomb logarithm $\ln\Lambda=3$ Eda et al. (2015). The mean rates of energy and momentum loss resulting from dynamical friction can be described as follows:
$\left(\frac{dE}{dt}\right)_{\text{DF}}$ $=$ $\bm{\bf f}_{\text{DF}}\cdot\bm{\bf v},$ (34)
$\left(\frac{d{\bf L}}{dt}\right)_{\text{DF}}$ $=$ $\bm{\bf r}\times\bm{\bf f}_{\text{DF}}.$ (35)
The averaged rates of energy and angular momentum loss due to GW emission are expressed as Peters and Mathews (1963)
$\left(\frac{dE}{dt}\right)_{\text{GW}}$ $=$ $-\frac{32}{5}\frac{\mu^{2}}{M^{2}}\left(\frac{M}{p}\right)^{5}f_{1}(e),$ (36)
$\left(\frac{dL}{dt}\right)_{\text{GW}}$ $=$ $-\frac{32}{5}\frac{\mu^{2}}{M^{2}}\left(\frac{M}{p}\right)^{7/2}f_{2}(e),$ (37)
@@
$f_{1}(e)$ $=$ $(1-e^{2})^{3/2}\left(1+\frac{73}{24}e^{2}+\frac{37}{96}e^{4}\right),$ (38)
$f_{2}(e)$ $=$ $(1-e^{2})^{3/2}\left(1+\frac{7}{8}e^{2}\right).$ (39)
@@
$\left(\frac{dE}{dt}\right)_{\text{orb}}$ $=$ $\left<\frac{dE}{dt}\right>_{\text{GW}}+\left<\frac{dE}{dt}\right>_{\text{acc}}+\left<\frac{dE}{dt}\right>_{\text{DF}},$ (40)
$\left(\frac{dL}{dt}\right)_{\text{orb}}$ $=$ $\left<\frac{dL}{dt}\right>_{\text{GW}}+\left<\frac{dL}{dt}\right>_{\text{DF}}.$ (41)
@@
$\left\langle\frac{dE}{dt}\right\rangle=\frac{1}{P}\int_{0}^{P}\frac{dE}{dt}\,dt=\frac{1}{P}\int_{0}^{2\pi}\frac{dE}{dt}\frac{dt}{d\phi}\,d\phi,$ (42)
$\left\langle\frac{dL}{dt}\right\rangle=\frac{1}{P}\int_{0}^{P}\frac{dL}{dt}\,dt=\frac{1}{P}\int_{0}^{2\pi}\frac{dL}{dt}\frac{dt}{d\phi}\,d\phi.$ (43)

### v2

@@ 1 Introduction
The interplay between black holes and DM has proved to be a significant field of study in theoretical astrophysics, especially in understanding the dynamics of gravitational systems [9, 10, 15]. Black holes play a crucial role in the formation and evolution of galaxies. On the other hand, DM, which constitutes about $27\%$ of the universe’s total mass-energy content, adds further complication to these gravitational systems and in cosmology [24]. In this context, the study of gravitational radiation reaction around black holes immersed in a DM halo offers deep insights into how these entities influence each other’s behavior and the resulting GW emitted.
@@
The presence of extended DM halos surrounding galaxies is strongly supported by various lines of evidence, including galaxy rotation curves, gravitational lensing analyses, and measurements of cosmic microwave background anisotropies. Among the models used to describe such halos, the Navarro–Frenk–White (NFW) profile [39], derived from cold DM simulations, is widely employed and is characterized by its steep inner density cusp. Nevertheless, the NFW profile has limited adaptability for representing alternative halo morphologies, such as cored density profiles. On the other hand, the Dehnen profile [23] offers a more general double power-law form, encompassing the Jaffe and Hernquist models as specific cases and featuring an adjustable inner slope. This versatility makes it especially suitable for capturing both cuspy and cored DM configurations, which are relevant for modeling the environments of black holes [45, 32]. It defines a one-parameter family of spherical equilibrium models in which the inner logarithmic density slope, $\gamma$, regulates the degree of central phase-space concentration. In contrast to fixed-shape prescriptions, such as the NFW profile, the Dehnen model provides a continuous interpolation between core-like ($\gamma\simeq 0$) and steeply cusped ($\gamma\simeq 2$) configurations without modifying its analytic form. This parametric freedom endows $\gamma$ with a physically interpretable role as an entropy-compression index, enabling the model to encapsulate a wide range of halo morphologies within a unified framework. Moreover, the model retains full analytic tractability in its cumulative mass, potential, and distribution function, rendering it particularly advantageous for dynamical modeling and gravitational-lensing applications while preserving a direct connection to the system’s inner information content.
@@
Recent advances in astrophysical research have increasingly highlighted the relevance of the Dehnen DM halo model. Since its introduction as a spherically symmetric density profile, the Dehnen model has gained widespread application in galactic dynamics and black hole astrophysics [23]. Initially employed to characterize the nuclear bulges of elliptical galaxies, it has more recently been used to construct black hole models embedded within Dehnen-profiled DM halos, revealing the impact of DM on the gravitational environment of black holes [40]. Further studies have derived black hole solutions for specific Dehnen profiles, systematically analyzing their thermodynamic behavior and strong-field gravitational signatures [29]. Investigations have also demonstrated that the gravitational lensing properties and orbital dynamics predicted by this model align well with general relativity, making it a preferred theory for explaining the peripheral environment of galaxies [5]. Collectively, these results establish the Dehnen profile as a powerful framework for probing DM–black hole interactions and the structure of galactic environments.
@@
In addition, GWs have become of large interest in gravitational physics since its detection by LIGO/Virgo collaboration in 2015 [1]. One of the main sources of GW is two body systems in which a stellar-mass compact object (SCO) rotates a supermassive black hole (SMBH). In such extreme mass ratio inspirals (EMRIs), low frequency GW would be produced which could be observed in future space-based observatories like LISA [6].
@@
In this paper, we are interested in EMRIs immersed in a DM halo whose density distribution is given by Dehnen type profile [22]. Dehnen distribution has been used, by several authors, to model the DM profile of dwarf galaxies [46, 34, 26]. This dark-matter background serves as a generalization of the well-established NFW profile, exhibiting a partially phenomenological character. In the following, we provide theoretical motivations for adopting this form. In addition, the traditional belief is that these galaxies rarely host a SMBH at their centers. However, some recent observations, including that of Leo I galaxy, show that SMBHs might be present at the center of some dwarf galaxies [11] (see also Ref. [43]). This has motivated study of black holes embedded in Dehnen type DM halo [41, 2, 30, 4, 3]. As the SCO spiral into the SMBH, it emits GWs. Further, the DM halo exerts a gravitational drag force on SCO which is known as dynamical friction. On the other hand, SCO accretes DM as it is moving inside the halo. We are interested in studying the evolution of the orbital parameter of SCO as well as its energy and angular momentum due to GW radiation reaction, dynamical friction, and the accretion. Various black hole solution surrounded by DM halo have been found in literature [18, 16, 17, 27] and the evolution of orbits of small compact objects inspiralling around these black holes has been investigated [21, 48]. These studies could help us find out about the existence and properties of DM.
@@
The paper is structured as follows. In Sec. 2, we describe the main features of dark matter density profile, computing the corresponding mass associated with dark matter. In Sec. 3, the bound orbits of the underlying metric is explored. In Sec. 4, the gravitational waveforms are studied, and finally in Sect. 5, we sketch our conclusions and perspectives.
2 Dark matter density profile and metric function
@@
$M_{D}=4\pi\int\limits_{0}^{r}\rho\left(r^{\prime}\right)r^{\prime 2}dr^{\prime}.$ (2.1)
@@
$\rho=\rho_{s}\left(\frac{r}{r_{s}}\right)^{-\gamma}\left[\left(\frac{r}{r_{s}}\right)^{\alpha}+1\right]^{\frac{\gamma-\beta}{\alpha}},$ (2.2)
where $\rho_{s}$ and $r_{s}$ denote the central halo density and the halo core radius respectively, whereas $\gamma$ determines the specific variant of the profile. The values of $\gamma$ lies within $[0,3]$, e.g. $\gamma=3/2$ is used to fit the surface brightness profiles of elliptical galaxies which closely resembles the de Vaucouleurs $r^{1/4}$ profile [44].
In general, the Dehnen model offers an improved representation of the dark-matter density profiles observed in $N$-body simulations. Owing to its analytic flexibility, the Dehnen family can accommodate physically motivated variations in the inner structure of halos. In particular, as demonstrated in Ref. [31], the adiabatic growth of a supermassive black hole (SMBH) at the halo center transforms an initial power-law cusp, $\rho_{i}(r)\propto r^{-\gamma}$ with $0<\gamma<2$, into a final density spike, $\rho_{f}(r)\propto r^{-\gamma_{\mathrm{sp}}}$, where $2.25<\gamma_{\mathrm{sp}}<2.5$. The adopted Dehnen profile with $\gamma=2.5$ thus reproduces the theoretically expected inner asymptotic behavior of such spiked halos within a fully analytic framework. By contrast, the standard NFW model–with its fixed inner slope $\gamma=1$–is incapable of capturing this physically motivated steep cusp. The Dehnen formulation, therefore, provides a more realistic and analytically tractable description of black hole-halo systems.
@@
Accordingly, Eq. (2.2) becomes
$\rho=\frac{\rho_{s}}{\left(\frac{r}{r_{s}}\right)^{5/2}\left(\frac{r}{r_{s}}+1\right)^{3/2}},$ (2.3)
and, so, plugging Eq. (2.3) in Eq. (2.1), one obtains
$M_{D}=4\pi\int\limits_{0}^{r}\frac{\rho_{s}r^{\prime 2}}{\left(\frac{r^{\prime}}{r_{s}}\right)^{5/2}\left(\frac{r^{\prime}}{r_{s}}+1\right)^{3/2}}dr^{\prime}=\frac{8\pi\rho_{s}r_{s}^{3}}{\sqrt{1+\frac{r_{s}}{r}}}.$ (2.4)
The static black hole metric in the Dehnen-$(1,4,5/2)$ DM halo is therefore given by  [^1] The authors of Ref. [2] model DM using an anisotropic fluid stress-energy tensor. Similarly, an effective anisotropic fluid has been proposed in Refs. [13, 12, 28] to unify dark energy and DM. The procedure to model, with an external energy-momentum tensor, induced by DM, a background metric appears a viable approximation to characterize with one metric only, two different spacetimes, i.e., an interior and an exterior configuration. Clearly, this limits the overall description, since it exhibits the presence of pressures violating the the cosmological cold DM hypothesis. However, it is possible to handle this technique at first approximation, albeit a more precise and complete scheme would require the Israel-Darmois junction conditions [20] or alternatives, see e.g. [36, 35].  [2]
$ds^{2}=-f\left(r\right)dt^{2}+\frac{dr^{2}}{f\left(r\right)}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (2.5)
@@
$f\left(r\right)=1-\frac{2M}{r}-32\pi\rho_{s}r_{s}^{3}\sqrt{\frac{r+r_{s}}{r_{s}^{2}\,r}}\,.$ (2.6)
This choice, though restrictive compared to the more general ansatz $g_{tt}=-e^{\Phi(r)}f(r)$, which can describe spacetimes with anisotropic matter distributions, is consistent with the weak-field origin of the DM halo model adopted in Ref. [2] and ensures compliance with standard energy conditions. Note that this geometry differs from the vacuum Schwarzschild black hole and the inclusion of DM has changed the horizon radius (see e.g. Ref. [47]).
3 Bound orbits of the corresponding black hole
@@ 3.1 The geodesic equations
$\frac{dt}{d\tau}=\frac{E}{f(r)},$ (3.1)
$\frac{d\phi}{d\tau}=\frac{L}{r^{2}},$ (3.2)
$\left(\frac{dr}{d\tau}\right)^{2}+V(r;L)=E^{2},$ (3.3)
where $\tau$ is the particle’s proper time [^2] Since the analysis is restricted to the weak-field region where $f(r)\approx 1$, the difference between $t$ and proper time $\tau$ is negligible. This is a standard simplification for large-radius orbits in similar studies.; $E$ and $L$ are constants of motion, respectively, the orbital energy and angular momentum; the effective potential for radial motion is given by
$V(r;L)=f(r)\left(1+\frac{L}{r^{2}}\right).$ (3.4)
@@ 3.2 Orbital parameters: $p$ and $e$
$\frac{\partial V}{\partial r}=\frac{\partial^{2}V}{\partial r^{2}}=0.$ (3.5)
@@
$L_{\rm bound}=\frac{2\sqrt{3}\left(M+8\pi\rho_{s}r_{s}^{3}\right)}{1-32\rho_{s}r_{s}^{2}}.$ (3.6)
@@
The roots of Eq. (3.7) can be obtained by defining [33]
$R_{3}(r)=r^{3}+\mathcal{A}_{1}r^{2}+\mathcal{A}_{2}r+\mathcal{A}_{3}\;,$ (3.8)
@@
$\mathcal{A}_{1}$ $=$ $-\frac{2M+16\pi\rho_{s}r_{s}^{3}}{(1-E^{2}-32\pi\rho_{s}r_{s}^{2})}\;,$ (3.9)
$\mathcal{A}_{2}$ $=$ $\frac{L^{2}(1-32\pi\rho_{s}r_{s}^{2})}{(1-E^{2}-32\pi\rho_{s}r_{s}^{2})}\;,$ (3.10)
$\mathcal{A}_{3}$ $=$ $-\frac{2L^{2}(M+8\pi\rho_{s}r_{s}^{3})}{(1-E^{2}-32\pi\rho_{s}r_{s}^{2})}\;.$ (3.11)
@@
$\mathcal{Q}$ $=$ $\frac{1}{9}\left(\mathcal{A}_{1}^{2}-3\mathcal{A}_{2}\right)\;,\;$ (3.12)
$\mathcal{R}$ $=$ $\frac{1}{54}\left(2\mathcal{A}_{1}^{3}-9\mathcal{A}_{1}\mathcal{A}_{2}+27\mathcal{A}_{3}\right),$ (3.13)
@@
$\vartheta=\arccos\left(\mathcal{R}/\sqrt{\mathcal{Q}^{3}}\right)\;.$ (3.14)
@@
$r_{\rm 1}$ $=$ $-2\sqrt{\mathcal{Q}}\cos\left(\frac{\vartheta+2\pi}{3}\right)-\frac{\mathcal{A}_{1}}{3}\;,$ (3.15)
$r_{\rm 2}$ $=$ $-2\sqrt{\mathcal{Q}}\cos\left(\frac{\vartheta-2\pi}{3}\right)-\frac{\mathcal{A}_{1}}{3}\;,$ (3.16)
$r_{3}$ $=$ $-2\sqrt{\mathcal{Q}}\cos\left(\frac{\vartheta}{3}\right)-\frac{\mathcal{A}_{1}}{3}\;.$ (3.17)
For a SCO inspiraling into a SMBH surrounded by a DM halo, the geodesic motion of the SCO can be parameterized by the semimajor axis $a$ and eccentricity $e$ through [7]
$r=\frac{a(1-e^{2})}{1+e\cos\chi},$ (3.18)
@@
For eccentric orbits, it is useful to relabel the outer two roots, which define the range of the smaller body’s motion: the outermost root named as the apoapsis, and the next root as the periapsis: $r_{1}=r_{a}$, $r_{2}=r_{p}$. From Eqs. (3.15) - (3.17), the third root can be written in terms of $p$ and $e$. Stable eccentric orbits occur when the inequality $r_{3}<r_{2}<r_{1}$ is satisfied, while stable circular orbits occur when $r_{1}=r_{2}>r_{3}$. The relations between $\left\{p,e\right\}$ and $\left\{E,L\right\}$ can be obtained as
$\epsilon$ $=$ $E/\mu=\sqrt{\frac{f(r_{a})f(r_{p})\left(r_{a}^{2}-r_{p}^{2}\right)}{r_{a}^{2}f(r_{p})-r_{p}^{2}f(r_{a})}},$ (3.19)
$\ell$ $=$ $L/\mu=\sqrt{\frac{r_{a}^{2}r_{p}^{2}\left(f(r_{a})-f(r_{p})\right)}{r_{a}^{2}f(r_{p})-r_{p}^{2}f(r_{a})}}.$ (3.20)
Combining Eq. (3.3) and Eqs. (3.18)-(3.20) yields
@@
$P$ $=$ $\int_{0}^{2\pi}\frac{dt}{d\chi}d\chi,$ (3.23)
$\Delta\phi$ $=$ $\int_{0}^{2\pi}\frac{d\phi}{d\chi}d\chi-2\pi.$ (3.24)
@@
EMRIs occurring within DM halos emit GWs and are affected by their ambient environment. The SCO’s motion is shaped by several forces, including GW backreaction, dynamical friction, and material accretion from the halo. If the SCO is a small black hole, it will draw in surrounding DM as it travels through the halo, a process that can be modeled using the Bondi-Hoyle accretion formula [37]
$\dot{\mu}=\frac{4\pi\mu^{2}\rho_{\text{DM}}}{(v^{2}+c_{\text{s}}^{2})^{3/2}},$ (3.25)
@@
$\left(\frac{dE}{dt}\right)_{\text{acc}}=\frac{\dot{\mu}}{\mu}E+\mu\dot{\epsilon}=\frac{\dot{\mu}}{\mu}\left(E-{\ell}\frac{dp}{d\ell}\frac{dE}{dp}\right).$ (3.26)
As the small black hole moves through a DM environment, it experiences gravitational interactions with DM particles within the halo. These interactions produce dynamical friction, which contributes to the loss of both energy and angular momentum. The Chandrasekhar formula for dynamical friction is expressed as [19]:
${\bf f}_{\text{DF}}=-\frac{4\pi\mu^{2}\rho_{\text{DM}}\ln\Lambda}{v^{3}}{\bf v},$ (3.27)
where the Coulomb logarithm $\ln\Lambda=3$ [25]. The mean rates of energy and momentum loss resulting from dynamical friction can be described as follows:
$\left(\frac{dE}{dt}\right)_{\text{DF}}$ $=$ ${\bf f}_{\text{DF}}\cdot{\bf v},$ (3.28)
$\left(\frac{d{\bf L}}{dt}\right)_{\text{DF}}$ $=$ ${\bf r}\times{\bf f}_{\text{DF}}.$ (3.29)
The averaged rates of energy and angular momentum loss due to GW emission are expressed as [42]
$\left(\frac{dE}{dt}\right)_{\text{GW}}$ $=$ $-\frac{32}{5}\frac{\mu^{2}}{M^{2}}\left(\frac{M}{p}\right)^{5}f_{1}(e),$ (3.30)
$\left(\frac{dL}{dt}\right)_{\text{GW}}$ $=$ $-\frac{32}{5}\frac{\mu^{2}}{M^{2}}\left(\frac{M}{p}\right)^{7/2}f_{2}(e),$ (3.31)
@@
$f_{1}(e)$ $=$ $(1-e^{2})^{3/2}\left(1+\frac{73}{24}e^{2}+\frac{37}{96}e^{4}\right),$ (3.32)
$f_{2}(e)$ $=$ $(1-e^{2})^{3/2}\left(1+\frac{7}{8}e^{2}\right).$ (3.33)
@@
$\left(\frac{dE}{dt}\right)_{\text{orb}}$ $=$ $\left<\frac{dE}{dt}\right>_{\text{GW}}+\left<\frac{dE}{dt}\right>_{\text{acc}}+\left<\frac{dE}{dt}\right>_{\text{DF}},$ (3.34)
$\left(\frac{dL}{dt}\right)_{\text{orb}}$ $=$ $\left<\frac{dL}{dt}\right>_{\text{GW}}+\left<\frac{dL}{dt}\right>_{\text{DF}}.$ (3.35)
@@
$\left\langle\frac{dE}{dt}\right\rangle=\frac{1}{P}\int_{0}^{P}\frac{dE}{dt}\,dt=\frac{1}{P}\int_{0}^{2\pi}\frac{dE}{dt}\frac{dt}{d\phi}\,d\phi,$ (3.36)
$\left\langle\frac{dL}{dt}\right\rangle=\frac{1}{P}\int_{0}^{P}\frac{dL}{dt}\,dt=\frac{1}{P}\int_{0}^{2\pi}\frac{dL}{dt}\frac{dt}{d\phi}\,d\phi.$ (3.37)